OpenStudy (anonymous):

Solve for the variable. a - 16 = -10

OpenStudy (anonymous):

add 16

OpenStudy (merchandize):

a=-10+16 a=6

OpenStudy (anonymous):

add the 16 over so what is -10 + 16 = 6

OpenStudy (anonymous):

add 16 to both sides to eliminate the 16 on the left hand side. this gives you a=-10+16 a=6